A)mary decides to purchasea house that cost $262,500.00 the bank requires a 10% downpayment and will provide a 15 year mortgage at annual interest rate of 8.0%. Calculate the mortgage loan amount, the monthly payment required to amortize the loan and the total interest paid(assume all the payments are made).use four 4 deciaml places then round dollar values to tthe nearest penny.
Mortgage payments are made each month(monthly)
b) create an amortization schedule for the first 4 months of Marys mortgage.
C) What is Marys approximate remaining balance after 59 payments have been made?

Answer by stanbon(48522)

A)mary decides to purchasea house that cost $262,500.00 the bank requires a 10% downpayment and will provide a 15 year mortgage at annual interest rate of 8.0%. Calculate the mortgage loan amount, the monthly payment required to amortize the loan and the total interest paid(assume all the payments are made).use four 4 deciaml places then round dollar values to tthe nearest penny.
Mortgage payments are made each month(monthly)
b) create an amortization schedule for the first 4 months of Marys mortgage.
C) What is Marys approximate remaining balance after 59 payments have been made?
------------
After the 10% down payment the mortgaged amount is $236,250
Formula for monthly regular payment =[Pi]/[1-(1+i)^-N]
where P=principal borrowed; N=# of payments; i=periodic interest rate
---------
monthly payment=[236,250*.08/12)]/[1-(1+0.08/12)^-(12*15)]
=[1575]/[0.6976039477]]
=2257.72
Total money paid in 15 years= 12*15*2257.72=$406,389.60
Total interest paid = 406,389.60-236,250=$170,139.60
---------
I haven't figured out the procedure on parts b and c yet.
I suggest you post this again so others may have a look
at it while I delve into b and c.
Cheers,
Stan H.